### v1

# Remote State Estimation over Unreliable Channels with Unreliable Feedback: Fundamental Limits

Abstract
This article is concerned with networked estimation in a system composed of a source that is observed by a sensor, a remote monitor that needs to estimate the state of the source in real time, and a communication channel that connects the source to the monitor. The source is a partially observable dynamical process, and the communication channel is a packet-erasure channel with feedback. Our main objective is to obtain the fundamental performance limits of the underlying networked system in the sense of a causal tradeoff between the packet rate and the mean square error when both forward and backward channels are unreliable. We characterize an optimal coding policy profile consisting of a scheduling policy for the encoder and an estimation policy for the decoder. We complement our theoretical results with a numerical analysis, and compare the performance limits of the networked system in different communication regimes.
Index Terms:
communication channels, feedback, optimal policies, packet loss, state estimators, unreliability.
I Introduction
Cyber-physical systems represent complex real-time systems that intricately integrate computational algorithms with dynamical processes, thereby facilitating seamless interactions between the digital and physical realms[1]. These systems harness computation, communication, and control to optimize efficiency, adaptability, and automation across diverse domains, including smart factories, smart healthcare, and smart transportation. A significant challenge within cyber-physical systems arises from their inherent dynamism, which requires continuous status updating to capture the latest information about the physical environment. As a step towards addressing this challenge, in this article, we develop a framework for networked estimation in a system composed of a source that is observed by a sensor, a remote monitor that needs to estimate the state of the source in real time, and a communication channel that connects the source to the monitor. The source is a partially observable dynamical process, and the communication channel is a packet-erasure channel with feedback. This framework is motivated by cyber-physical systems operating over digital networks, where the quality of decision making tightly depends on the availability of sensory observations transmitted as data packets, while packet loss appears as an inevitable phenomenon due to the real-time nature of the operation. Our main objective in this article is to obtain the fundamental performance limits of the underlying networked system in the sense of a causal tradeoff between the packet rate and the mean square error when both forward and backward channels are unreliable. To that end, we seek for an optimal coding policy profile that determines the policies of an encoder and a decoder, which are collocated with the source and the monitor, respectively.
I-A Related Work
It is known that reliable communication close to the capacity limit can be attained with error correction subject to infinite delay or with persistent retransmission based on feedback. In networked estimation, where data is real-time, any delay more than a certain threshold is typically intolerable. Moreover, retransmission of stale information is unfavourable when fresh information can be transmitted with the same success rate instead. For these reasons, networked estimation is often conducted in an unreliable communication regime. Previous research has already recognized the severe effects of packet loss, as a consequence of communication unreliability, on stability of Kalman filtering [2, 3, 4, 5, 6, 7, 8, 9, 10]. In a seminal work, Sinopoli et al. [2] studied mean-square stability of Kalman filtering over an independent and identically distributed (i.i.d.) channel, and proved that there exists a critical point for the packet error rate above which the expected estimation error covariance is unbounded. To capture the temporal correlation of wireless channels, Wu et al. [6] addressed peak-covariance stability of Kalman filtering over a Gilbert-Elliott channel, and proved that there exists a critical region defined by the recovery and failure rates outside which the expected prediction error covariance is unbounded. In addition, to take into account the time variation of the strengths of wireless channels, Quevedo et al. [9] investigated mean-square stability of Kalman filtering over a fading channel with correlated gains, and established a sufficient condition that ensures the exponential boundedness of the expected estimation error covariance. Note that, the above works assume that the outputs of the source are transmitted in a periodic manner, and seek to derive fundamental stability conditions regardless of the existence of feedback.
To determine the fundamental performance limits of networked estimation, however, one requires to make a tradeoff between communication and estimation costs. Various classes of such a tradeoff have been explored in the literature [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. In particular, the tradeoff between the packet rate and the mean square error, in which a signal-dependent scheduling policy and an estimation policy need to be found, was considered in [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 30, 31]. In this class, which is of main interest to our study, Lipsa and Martins [12, 13] analysed the estimation of a scalar Gauss–Markov process over a lossless channel and an i.i.d. channel with reliable feedback, and proved based on majorisation theory that the optimal scheduling policy is symmetric threshold and the optimal estimation policy is linear. Molin and Hirche [14] investigated the estimation of a scalar Markov process with symmetric noise distribution over a lossless channel, and devised an iterative algorithm that produces policies similar to those found in [12]. Chakravorty and Mahajan [15, 16, 17] studied the estimation of a scalar autoregressive Markov process with symmetric noise distribution over a lossless channel, an i.i.d. channel with reliable feedback, as well as a Gilbert–Elliott channel with reliable feedback, and proved based on renewal theory that the optimal scheduling policy is symmetric threshold and the optimal estimation policy is linear. Moreover, Guo and Kostina [19] addressed the estimation of the scalar Ornstein–Uhlenbeck process over a lossless channel, and based on majorisation theory and real induction proved that the optimal scheduling policy remains symmetric threshold and the optimal estimation policy remains linear.
Note that, in networked estimation, feedback can serve as a mechanism that helps the encoder know the belief of the decoder about the states of the source. Previous works on networked estimation have mostly assumed that feedback is perfect. However, when feedback is imperfect, the knowledge of the encoder about the belief of the decoder becomes distorted. This consequently can cause a degradation in the system performance. Despite its critical importance, there has been limited research on networked estimation with unreliable feedback [32, 33, 34, 35, 36, 37]. Notably, Nourian et al. [32, 33] studied the estimation of a Gauss–Markov process over a lossy channel with erroneous feedback. The goal in [32] is to minimize the trace of error variance subject to an energy harvesting constraint, and that in [33] is to minimize a convex combination of the trace of error variance and the communication energy. The authors of these articles used the notion of information state to transform the problem into a simplified one, and accordingly derived the optimal signal-independent scheduling policy and the optimal estimation policy. Khina et al. [34] studied the sequential coding of a Gauss–Markov process with multiple sensors over an i.i.d. channel with delayed feedback, and characterized the achievable causal rate-distortion region. Furthermore, Guo et al. [35] and Li et al. [36, 37] studied the estimation of a Gauss–Markov process over a lossy channel with compromised feedback, and, for some fixed scheduling policies, derived attack policies that target packet acknowledgment signals. The above results show that erroneous, delayed, or compromised feedback in general degrades the system performance. Note that, differently from the focus of our study, feedback can also be used for increasing the channel reliability, provisioning of the channel states, or signifying the receipts of control commands. For the effects of feedback in these cases, we refer the reader to [38, 39, 40, 41, 42], to [43, 44, 45, 46], and to [47, 48, 49, 50, 51, 52, 53, 54, 55], respectively.
I-B Contributions and Outline
In this article, we establish a theoretical framework for networked estimation over unreliable channels with unreliable feedback. We characterize a globally optimal coding policy profile consisting of a symmetric threshold scheduling policy for the encoder and a linear estimation policy for the decoder. We show that the encoder requires to run four recursive equations, while the decoder requires to run only one recursive equation. We complement our theoretical results with a numerical analysis, and compare the performance limits of the networked system in different communication regimes. A key observation is about the variations of the system performance with respect to the feedback error rate.
We emphasise that our findings present innovative contributions to the field of goal-oriented communication [56]. Our study deviates drastically from prior works. More specifically, the studies in [2, 3, 4, 5, 6, 7, 8, 9, 10] deal with networked estimation stability. In contrast, our results focus on networked estimation performance. The studies in [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21] consider scalar sources and lossless feedback. On the contrary, our results hold for multi-dimensional sources and lossy feedback. Finally, the studies in [22, 23, 24, 32, 33, 35, 36, 37] derive variance-based scheduling policies that do not take advantage of realized measurements, and the studies in [25, 26, 27, 28, 29] derive compression policies that operate in a periodic manner. Differently, our results characterize a scheduling policy that depends on realized measurements and operates in a non-periodic manner.
The article is organized in five sections. The current introductory section is followed by Section II, where we formulate a causal tradeoff. We present and discuss our main theoretical result in Sections III, and provide our numerical analysis in Section IV. Finally, we conclude the article in Section V.
II Problem Statement
In this section, we mathematically formulate the problem of interest. In our networked system, the state of the source is observed by a sensor. The channel, connecting the sensor to the monitor, is composed of a forward and a backward channel. The encoder, collocated with the source, and the decoder, collocated with the monitor, constitute the two distributed decision makers of the networked system. We first introduce the models of the source and the channel, and then formulate a causal tradeoff between the packet rate and the mean square error. In our models, we consider a finite time horizon $N$, and assume that the time is discretized into time slots with constant durations.
II-A Source Model
The source is partially observable Gauss–Markov process, and is described by the discrete-time state and output equations
$x_{k+1}$ $=A_{k}x_{k}+w_{k}$ (1)
$y_{k}$ $=C_{k}x_{k}+v_{k}$ (2)
for $k\in\mathbb{N}_{[0,N]}$ with the initial condition $x_{0}$, where $x_{k}\in\mathbb{R}^{n}$ is the state of the process, $A_{k}\in\mathbb{R}^{n\times n}$ is the state matrix, $w_{k}\in\mathbb{R}^{n}$ is a Gaussian white noise with zero mean and covariance $W_{k}\succ 0$, $y_{k}\in\mathbb{R}^{m}$ is the output of the process, $C_{k}\in\mathbb{R}^{m\times n}$ is the output matrix, and $v_{k}\in\mathbb{R}^{m}$ is a Gaussian white noise with zero mean and covariance $V_{k}\succ 0$.
Assumption 1
For the source model in (1) and (2), the following assumptions are satisfied:
  1. (i)
    The initial condition $x_{0}$ is a Gaussian vector with mean $m_{0}$ and covariance $M_{0}$.
  2. (ii)
    The random variables $x_{0}$, $w_{t}$, and $v_{s}$ for $t,s\in\mathbb{N}_{[0,N]}$ are mutually independent, i.e., $\operatorname{\mathsf{p}}(x_{0},\boldsymbol{w}_{0:N},\boldsymbol{v}_{0:N})=\operatorname{\mathsf{p}}(x_{0})\prod_{k=0}^{N}\operatorname{\mathsf{p}}(w_{k})\prod_{k=0}^{N}\operatorname{\mathsf{p}}(v_{k})$.
Remark 1
The source model given in (1) and (2) represents a wide class of physical processes. Note that such a Gauss–Markov model has been adopted extensively in the fields of communication and control, as studying this model can provide a foundation for development of more sophisticated systems. Moreover, the time-varying nature of the model allows approximation of nonlinear systems around their nominal trajectories, and the partially observable nature of it respects the fact that in reality only a noisy version of the output can be observed. Clearly, a model that is time-invariant and/or fully observable is a special case of this model.
II-B Channel Model
The communication channel is a packet-erasure channel with feedback. At each time $k$, a message containing sensory information, represented by $\check{x}_{k}$, can be sent by the encoder over the forward channel, which has one-step delay and is i.i.d with forward error rate $\lambda$, to the decoder, where an estimate of the state of the process, represented by $\hat{x}_{k}$, should be computed. Let $\delta_{k}\in\{0,1\}$ be a binary variable such that $\delta_{k}=1$ if $\check{x}_{k}$ is transmitted at time $k$, and $\delta_{k}=0$ otherwise. A transmitted message in the forward channel at time $k$ is successfully received by the decoder with probability $\lambda^{c}:=1-\lambda$. Let $\gamma_{k}\in\{0,1\}$ be a binary random variable modelling the packet loss in the forward channel such that $\gamma_{k}=0$ if a packet loss occurs in the forward channel when it is used to exchange $\check{x}_{k}$, and $\gamma_{k}=1$ otherwise. Then, $\operatorname{\mathsf{p}}(\gamma_{k}=0)=\lambda$ for $k\in\mathbb{N}_{[0,N]}$. Accordingly, the output of the forward channel at time $k+1$ is specified as
$z^{f}_{k+1}=\left\{\begin{array}[]{l l}\check{x}_{k},&\ \text{if}\ \delta_{k}=1\ \wedge\ \gamma_{k}=1,\\[3.0pt] \mathfrak{F},&\ \text{if}\ \delta_{k}=1\ \wedge\ \gamma_{k}=0,\\[3.0pt] \mathfrak{E},&\ \text{otherwise}\end{array}\right.$ (6)
for $k\in\mathbb{N}_{[0,N]}$ with $z^{f}_{0}=\mathfrak{E}$ by convention, where $\mathfrak{F}$ and $\mathfrak{E}$ represent packet loss and absence of transmission in the forward channel, respectively. At each time $k$, if the forward channel is used to exchange $\check{x}_{k}$, i.e., if $\delta_{k}=1$, a packet acknowledgment is sent back by the decoder to the encoder over the backward channel, which has zero delay and is i.i.d. with backward error rate $\rho$. A transmitted message in the backward channel at time $k$ is successfully received by the encoder with probability $\rho^{c}:=1-\rho$. Let $\theta_{k}\in\{0,1\}$ be a binary random variable modelling the packet loss in the backward channel such that $\theta_{k}=0$ if a packet loss occurs in the backward channel when it is used to exchange $\gamma_{k}$, and $\theta_{k}=1$ otherwise. Then, $\operatorname{\mathsf{p}}(\theta_{k}=0)=\rho$ for $k\in\mathbb{N}_{[0,N]}$. Accordingly, the output of the backward channel at time $k+1$ is specified as
$z^{b}_{k+1}=\left\{\begin{array}[]{l l}\gamma_{k},&\ \text{if}\ \delta_{k}=1\ \wedge\ \theta_{k}=1,\\[3.0pt] \mathfrak{H},&\ \text{if}\ \delta_{k}=1\ \wedge\ \theta_{k}=0,\\[3.0pt] \mathfrak{G},&\ \text{otherwise}\end{array}\right.$ (10)
for $k\in\mathbb{N}_{[0,N]}$ with $z^{b}_{0}=\mathfrak{G}$ by convention, where $\mathfrak{H}$ and $\mathfrak{G}$ represent packet loss and absence of transmission in the backward channel, respectively.
Assumption 2
For the channel model in (6) and (10), the following assumptions are satisfied:
  1. (i)
    The random variables $\gamma_{t}$ and $\theta_{s}$ for $t,s\in\mathbb{N}_{[0,N]}$ are mutually independent, i.e., $\operatorname{\mathsf{p}}(\boldsymbol{\gamma}_{0:N},\boldsymbol{\theta}_{0:N})=\prod_{k=0}^{N}\operatorname{\mathsf{p}}(\gamma_{k})\prod_{k=0}^{N}\operatorname{\mathsf{p}}(\theta_{k})$.
  2. (ii)
    Quantization error is negligible, i.e., in a successful transmission, real-value sensory information can be conveyed from the encoder to the decoder without any error.
Remark 2
The channel model given in (6) and (10) lends itself to the representation of the different communication regimes. In particular, we have reliable communication when $\lambda=0$ and $\rho=1$, have unreliable communication with reliable acknowledgments when $\lambda>0$ and $\rho=1$, have unreliable communication without acknowledgments when $\lambda>0$ and $\rho=0$, and have unreliable communication with unreliable acknowledgments when $\lambda>0$ and $\rho>0$. Note that the forward and backward channels operate with one-step delay and zero delay, respectively, as the main communication load is in the forward channel. Moreover, the effect of the quantization error is negligible because the number of bits in each data packet is assumed to be sufficiently large, which is a common assumption in the literature of networked estimation [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24].
Remark 3
According to (6), aside from the sensory information that can be encoded explicitly at time $k$ as $\check{x}_{k}$, implicit information can be encoded through $\mathfrak{F}$ and $\mathfrak{E}$. To illustrate this, suppose that $\delta_{k}=1$ if and only if $y_{k}\in\mathcal{A}_{k}$, where $\mathcal{A}_{k}$ can be a measurable set. Then, the decoder can infer that $y_{k}\in\mathcal{A}_{k}$ or $y_{k}\notin\mathcal{A}_{k}$ when $\mathfrak{F}$ or $\mathfrak{E}$ is detected at time $k+1$, respectively. Besides, according to (10), $\mathfrak{H}$ is detected by the encoder at time $k+1$ when a packet loss occurs in the backward channel at time $k+1$, and $\mathfrak{G}$ is detected at time $k+1$ when nothing is transmitted in the forward channel at time $k$.
II-C A Team Decision-Making Problem
In our framework, the decision variables $\delta_{k}$ and $\hat{x}_{k}$ at each time $k$ for $k\in\mathbb{N}_{[0,N]}$ are determined by the encoder and the decoder in a distributed way, respectively. This distributed decision-making process relies on causal information. Let the information sets of the encoder and the decoder be denoted by $\mathcal{I}^{e}_{k}$ and $\mathcal{I}^{d}_{k}$. A coding policy profile $(\epsilon,\delta)$ consisting of a scheduling policy $\epsilon$ and an estimation policy $\delta$ is admissible if $\epsilon=\{\operatorname{\mathsf{P}}(\delta_{k}|\mathcal{I}^{e}_{k})\}_{k=0}^{N}$ and $\delta=\{\operatorname{\mathsf{P}}(\hat{x}_{k}|\mathcal{I}^{d}_{k})\}_{k=0}^{N}$, where $\operatorname{\mathsf{P}}(\delta_{k}|\mathcal{I}^{e}_{k})$ and $\operatorname{\mathsf{P}}(\hat{x}_{k}|\mathcal{I}^{d}_{k})$ are measurable stochastic kernels. In the sequel, we seek to find a globally optimal solution $(\epsilon^{\star},\delta^{\star})$ to a team decision-making problem cast as
$\underset{\epsilon\in\mathcal{E},\delta\in\mathcal{D}}{\operatorname{minimize}}\ \Phi(\epsilon,\delta)$ (11)
subject to the source model in (1) and (2), and the channel model in (6) and (10), where $\mathcal{E}$ and $\mathcal{D}$ are the sets of admissible scheduling policies and admissible estimation policies, and
$\Phi(\epsilon,\delta):=\operatorname{\mathsf{E}}\bigg{[}\sum_{k=0}^{N}\alpha_{k}\delta_{k}+(x_{k}-\hat{x}_{k})^{T}(x_{k}-\hat{x}_{k})\bigg{]}$ (12)
for the weighting coefficient $\alpha_{k}\geq 0$, which specifies communication cost per transmission. The concept of global optimality in the context of our team decision-making problem is captured by the following definition.
Definition 1 (Global optimality)
A policy profile $(\epsilon^{\star},\delta^{\star})$ in the above team decision-making problem with two distributed decision makers is globally optimal if
$\Phi(\epsilon^{\star},\delta^{\star})\leq\Phi(\epsilon,\delta),\ \text{for all }\epsilon\in\mathcal{E},\delta\in\mathcal{D}.$
Remark 4
The loss function in (12) is expressed in terms of two performance criteria: the mean-square-error and the packet-rate [^1] In the loss function in (12), the coefficient $1/N$ that appears in the definitions of these performance criteria is omitted, as it has no effect on the solutions.. The mean-square-error criterion, which is frequently used in the analysis of control systems, represents the quality of estimation. However, the packet-rate criterion, which is frequently used in the analysis of packet-switching networks, represents the cost of communication. We should emphasise that the problem (11) is a decision-making problem with a non-classical information structure. This problem is nonconvex and in general intractable. Finding a globally optimal solution $(\epsilon^{\star},\delta^{\star})$ to this problem allows us to determine the fundamental performance limits of the networked system in different communication regimes.
III Main Result
In this section, we present the main theoretical result on the synthesis of a globally optimal policy profile. Let $\check{x}_{k}=\operatorname{\mathsf{E}}[x_{k}|\mathcal{I}^{e}_{k}]$ and $\hat{x}_{k}=\operatorname{\mathsf{E}}[x_{k}|\mathcal{I}^{d}_{k}]$. In addition, define the innovation at the encoder $\nu_{k}:=y_{k}-C_{k}\operatorname{\mathsf{E}}[x_{k}|\mathcal{I}^{e}_{k-1}]$, the estimation error at the decoder $\hat{e}_{k}:=x_{k}-\operatorname{\mathsf{E}}[x_{k}|\mathcal{I}^{d}_{k}]$, the estimation mismatch $\tilde{e}_{k}:=\operatorname{\mathsf{E}}[x_{k}|\mathcal{I}^{e}_{k}]-\operatorname{\mathsf{E}}[x_{k}|\mathcal{I}^{d}_{k}]$, the conditional mean of the estimation mismatch at the encoder $\breve{e}_{k}:=\operatorname{\mathsf{E}}[\tilde{e}_{k}|\mathcal{I}^{e}_{k}]$, and the conditional covariance of the estimation mismatch at the encoder $R_{k}:=\operatorname{\mathsf{cov}}[\tilde{e}_{k}|\mathcal{I}^{e}_{k}]$.
In general, the information sets of the encoder and the decoder are specified as
$\mathcal{I}^{e}_{k}$ =$=\Big{\{}y_{t},z^{b}_{t},\delta_{s},\theta_{s}\big{|}t\in\mathbb{N}_{[0,k]},s\in\mathbb{N}_{[0,k-1]}\Big{\}}$
$\mathcal{I}^{d}_{k}$ =$=\Big{\{}z^{f}_{t},\delta_{s},\gamma_{s}\big{|}t\in\mathbb{N}_{[0,k]},s\in\mathbb{N}_{[0,k-1]}\Big{\}}$
The following definition introduces a value function associated with the loss function in (12).
Definition 2 (Value function)
The value function $V_{k}(\mathcal{I}^{e}_{k})$ associated with the loss function $\Phi(\epsilon,\delta)$ and the information set $\mathcal{I}^{e}_{k}$ is defined by
$V_{k}(\mathcal{I}^{e}_{k}):=$ $\min_{\epsilon\in\mathcal{E}:\delta=\delta^{\star}}\operatorname{\mathsf{E}}\bigg{[}\sum_{t=k}^{N-1}\alpha_{t}\delta_{t}+\hat{e}_{t+1}^{T}\hat{e}_{t+1}\Big{|}\mathcal{I}^{e}_{k}\bigg{]}$ (13)
for $k\in\mathbb{N}_{[0,N]}$, where the estimation policy $\delta^{\star}$ is chosen based on $\hat{x}_{k}=\operatorname{\mathsf{E}}[x_{k}|\mathcal{I}^{d}_{k}]$.
The value function $V_{k}(\mathcal{I}^{e}_{k})$ is a quantity defined from the perspective of the encoder, taking into account the fact that, due to causality, the scheduling policy $\epsilon\in\mathcal{E}$ at each time $k$ can affect $\hat{e}_{t}$ only from time $t=k+1$ onward.
The following theorem provides a globally optimal policy profile for the problem in (11).
Theorem 1
The tradeoff between the packet rate and the mean square error admits a globally optimal solution $(\epsilon^{\star},\delta^{\star})$ such that $\epsilon^{\star}$ is a symmetric threshold scheduling policy given by
$\delta_{k}=\mathds{1}_{\chi_{k}-\alpha_{k}\geq 0}$ (14)
in conjunction with $\check{x}_{k}=\operatorname{\mathsf{E}}[x_{k}|\mathcal{I}^{e}_{k}]$ for $k\in\mathbb{N}_{[1,N]}$, where $\chi_{k}(\breve{e}_{k},R_{k})=\lambda^{c}\breve{e}_{k}^{T}A_{k}^{T}A_{k}\breve{e}_{k}+\lambda^{c}\operatorname{tr}(A_{k}R_{k}A_{k}^{T})+\operatorname{\mathsf{E}}[V_{k+1}(\mathcal{I}^{e}_{k+1})|\mathcal{I}^{e}_{k},\delta_{k}=0]-\operatorname{\mathsf{E}}[V_{k+1}(\mathcal{I}^{e}_{k+1})|\mathcal{I}^{e}_{k},\delta_{k}=1]$ is a symmetric function of $\breve{e}_{k}$ and a function $R_{k}$, which requires solving
$\check{x}_{k}$ $=A_{k-1}\check{x}_{k-1}+K_{k}\nu_{k}$ (15)
$Q_{k}$ $=\big{(}(A_{k-1}Q_{k-1}A_{k-1}^{T}+W_{k-1})^{-1}+L_{k}\big{)}^{-1}$ (16)
$\breve{e}_{k}$ $=(1-\bar{\theta}_{k-1})A_{k-1}\breve{e}_{k-1}+K_{k}\nu_{k}$ (17)
$R_{k}$ $=(1-\bar{\theta}_{k-1})A_{k-1}R_{k-1}A_{k-1}^{T}+\bar{\bar{\theta}}_{k-1}H_{k-1}$ (18)
for $k\in\mathbb{N}_{[1,N]}$ with initial conditions $\check{x}_{0}=m_{0}+K_{0}\nu_{0}$, $Q_{0}=(M_{0}^{-1}+C_{0}^{T}V_{0}^{-1}C_{0})^{-1}$, $\breve{e}_{0}=K_{0}\nu_{0}$, and $R_{0}=0$, where $K_{k}=Q_{k}C_{k}^{T}V_{k}^{-1}$, $L_{k}=C_{k}^{T}V_{k}^{-1}C_{k}$, $\bar{\theta}_{k-1}=\theta_{k-1}\delta_{k-1}\gamma_{k-1}+\lambda^{c}(1-\theta_{k-1})\delta_{k-1}$, $\bar{\bar{\theta}}_{k-1}=(1-\theta_{k-1})(\lambda^{c}\delta_{k-1}-{\lambda^{c}}^{2}\delta_{k-1})$, and $H_{k-1}=A_{k-1}\breve{e}_{k-1}\breve{e}_{k-1}^{T}A_{k-1}^{T}$; and $\delta^{\star}$ is a non-Gaussian linear estimation policy given by
$\hat{x}_{k}=A_{k-1}\hat{x}_{k-1}+\delta_{k-1}\gamma_{k-1}A_{k-1}(\check{x}_{k-1}-\hat{x}_{k-1})$ (19)
for $k\in\mathbb{N}_{[1,N]}$ with initial condition $\hat{x}_{0}=m_{0}$, where $\hat{x}_{k}=\operatorname{\mathsf{E}}[x_{k}|\mathcal{I}^{d}_{k}]$, without being influenced by any signaling effect.
Proof:
See the Appendix. ∎
Remark 5
In the structural results of Theorem 1, the scheduling policy $\epsilon^{\star}$ at each time $k$ depends on $\breve{e}_{k}$ and $R_{k}$, which satisfy the recursive equations (17) and (18). The preceding terms in (17) and (18) can be eliminated only if $\delta_{k-1}\gamma_{k-1}\theta_{k-1}=1$. In addition, the estimation policy $\delta^{\star}$ at each time $k$ is given by $\hat{x}_{k}$, which satisfies the recursive equation (19). The preceding term in (19) is eliminated only if $\delta_{k-1}\gamma_{k-1}=1$. Note that the complexity of computing the policy profile $(\epsilon^{\star},\delta^{\star})$ is $\mathcal{O}(Nd_{0}d_{1}^{n}d_{2}^{n(n+1)/2})$ when the expected value is obtained based on a weighted sum of $d_{0}$ samples, and $\breve{e}_{k}$ and $R_{k}$ are discretized in grids with $d_{1}^{n}$ and $d_{2}^{n(n+1)/2}$ points, respectively.
IV Numerical Analysis
In this section, we provide a numerical analysis to demonstrate and compare the fundamental performance limits of the studied networked system in different communication regimes. Consider the following parameters: the state coefficient $A_{k}=0.9$, the output coefficient $C_{k}=1$, the process noise variance $W_{k}=3$, the sensor noise variance $V_{k}=1$ for $k\in\mathbb{N}_{[0,N]}$, the mean and the variance of the initial state $m_{0}=0$ and $M_{0}=1$, and the time horizon $N=100$. Assume that the forward error rate $\lambda$, the backward error rate $\rho$, and the weighting coefficient $\alpha_{k}$ remain constant over the whole time horizon.
We used the Monte Carlo simulation and obtained our results for different values of the forward error rate $\lambda$, the backward error rate $\rho$, and the weighting coefficient $\alpha$; and assessed the impacts of feedback on the fundamental performance limits. Figs. 1 and 2 depict the causal tradeoff curves when the forward error rate $\lambda$ is $0.1$ and $0.3$, respectively, and the backward error rate $\rho$ takes values in $\{0.0,0.3,0.7,1.0\}$. Note that, in these diagrams, the regions above curves are achievable. This implies that there exists no coding policy profile with performance outside these achievable regions. As expected, we observe that the lower the backward error rate, the better the system performance.
Figure 1: The causal tradeoff curves when $\lambda=0.1$ and $\rho\in\{0.0,0.3,0.7,1.0\}$.
Figure 2: The causal tradeoff curves when $\lambda=0.3$ and $\rho\in\{0.0,0.3,0.7,1.0\}$.
V Conclusions
In this article, we established a theoretical framework for networked estimation over unreliable channels with unreliable feedback. We characterized a globally optimal coding policy profile consisting of a symmetric threshold scheduling policy for the encoder and a linear estimation policy for the decoder. In addition, we obtained the recursive equations that need to be solved by the encoder and the decoder, and specified the computational complexity required for the implementation of the optimal policies. Finally, we provided a numerical analysis to demonstrate and compare the fundamental performance limits of the networked system in different communication regimes.
Appendix: Derivation of Main Results
This section is dedicated to the derivation of the main results. The following Lemma shows that the conditional means $\operatorname{\mathsf{E}}[x_{k}|\mathcal{I}^{e}_{k}]$ and $\operatorname{\mathsf{E}}[x_{k}|\mathcal{I}^{d}_{k}]$ are instrumental for the operations performed by the encoder and the decoder.
Lemma 1
Without loss of optimality, at each time $k$, one can adopt $\check{x}_{k}=\operatorname{\mathsf{E}}[x_{k}|\mathcal{I}^{e}_{k}]$ as the message that can be transmitted by the encoder, and $\hat{x}_{k}=\operatorname{\mathsf{E}}[x_{k}|\mathcal{I}^{d}_{k}]$ as the state estimate that can be computed by the decoder.
Proof:
Let $\epsilon^{o}$ be a globally optimal scheduling policy that is implemented. Given all information available at the decoder, $\operatorname{\mathsf{E}}[x_{k}|\mathcal{I}^{d}_{k}]$ is the optimal value that minimizes the mean square error at the decoder at time $k$. Moreover, $\operatorname{\mathsf{E}}[x_{k}|\mathcal{I}^{e}_{k}]$ fuses all the previous and current outputs of the process that are available at the encoder at time $k$. This implies that from the minimum mean-square-error perspective the decoder is able to develop a state estimate upon the receipt of a message at time $k+1$, that would be the same if it had all the previous outputs of the process until time $k$, which is the best possible case for the decoder. Hence, both claims hold. ∎
The next two lemmas characterize the optimal estimators at the encoder and the decoder.
Lemma 2
The optimal estimator minimizing the mean square error at the encoder satisfies the recursive equations
$\check{x}_{k}$ $=m_{k}+K_{k}(y_{k}-C_{k}m_{k})$ (20)
$m_{k}$ $=A_{k-1}\check{x}_{k-1}$ (21)
$Q_{k}$ $=(M_{k}^{-1}+C_{k}^{T}V_{k}^{-1}C_{k})^{-1}$ (22)
$M_{k}$ $=A_{k-1}Q_{k-1}A_{k-1}^{T}+W_{k-1}$ (23)
for $k\in\mathbb{N}_{[1,N]}$ with initial conditions $\check{x}_{0}=m_{0}+K_{0}(y_{0}-C_{0}m_{0})$ and $Q_{0}=(M_{0}^{-1}+C_{0}^{T}V_{0}^{-1}C_{0})^{-1}$, where $m_{k}=\operatorname{\mathsf{E}}[x_{k}|\mathcal{I}^{e}_{k-1}]$, $K_{k}=Q_{k}C_{k}^{T}V_{k}^{-1}$, $Q_{k}=\operatorname{\mathsf{cov}}[x_{k}|\mathcal{I}^{e}_{k}]$, and $M_{k}=\operatorname{\mathsf{cov}}[x_{k}|\mathcal{I}^{e}_{k-1}]$.
Proof:
Observe that, given $\mathcal{I}^{e}_{k}$, the minimum mean-square-error estimator is $\operatorname{\mathsf{E}}[x_{k}|\mathcal{I}^{e}_{k}]$. It is evident that this estimator satisfies the Kalman filter equations (see, e.g., [57]). ∎
Lemma 3
The optimal estimator minimizing the mean square error at the decoder satisfies the recursive equation
$\hat{x}_{k}$ $=A_{k-1}\hat{x}_{k-1}+\delta_{k-1}\gamma_{k-1}A_{k-1}(\check{x}_{k-1}-\hat{x}_{k-1})$
$\quad+(1-\delta_{k-1})\imath_{k-1}+\delta_{k-1}(1-\gamma_{k-1})\jmath_{k-1}$ (24)
for $k\in\mathbb{N}_{[1,N]}$ with initial condition $\hat{x}_{0}=m_{0}$, where $\imath_{k-1}=A_{k-1}\operatorname{\mathsf{E}}[\hat{e}_{k-1}|\mathcal{I}_{k-1}^{d},\delta_{k-1}=0]$ and $\jmath_{k-1}=A_{k-1}\operatorname{\mathsf{E}}[\hat{e}_{k-1}|\mathcal{I}_{k-1}^{d},\delta_{k-1}=1]$.
Proof:
Observe that, given $\mathcal{I}^{d}_{k}$, the minimum mean-square-error estimator is $\operatorname{\mathsf{E}}[x_{k}|\mathcal{I}^{d}_{k}]$. Taking the conditional expectation of (1) given $\mathcal{I}^{d}_{k}$, we obtain
$\operatorname{\mathsf{E}}\Big{[}x_{k}\big{|}\mathcal{I}^{d}_{k}\Big{]}=A_{k-1}\operatorname{\mathsf{E}}\Big{[}x_{k-1}\big{|}\mathcal{I}_{k}^{d}\Big{]}$ (25)
for $k\in\mathbb{N}_{[1,N]}$ as $w_{k-1}$ is independent of $\mathcal{I}^{d}_{k}$ and has zero mean. When sensory information is successfully delivered in the forward channel at time $k$, we have $z^{f}_{k}=\check{x}_{k-1}$. In this case, we get $\operatorname{\mathsf{E}}[x_{k-1}|\mathcal{I}^{d}_{k}]=\operatorname{\mathsf{E}}[x_{k-1}|\check{x}_{k-1},Q_{k-1}]=\check{x}_{k-1}$ as $\{\check{x}_{k-1},Q_{k-1}\}$ is a sufficient statistic of $\mathcal{I}^{e}_{k}$. Hence, using (25), when $z^{f}_{k}=\check{x}_{k-1}$, i.e., when $\delta_{k-1}=1\wedge\gamma_{k-1}=1$, we get
$\operatorname{\mathsf{E}}\Big{[}x_{k}\big{|}\mathcal{I}^{d}_{k}\Big{]}=A_{k-1}\check{x}_{k-1}$ (26)
for $k\in\mathbb{N}_{[1,N]}$. However, when no sensory information is delivered in the forward channel at time $k$, we have $z^{f}_{k}=\mathfrak{E}$ or $z^{f}_{k}=\mathfrak{F}$. Define $p_{k-1}:=\operatorname{\mathsf{E}}[x_{k-1}|\mathcal{I}^{d}_{k-1},\delta_{k-1}=0]-\operatorname{\mathsf{E}}[x_{k-1}|\mathcal{I}^{d}_{k-1}]$ when $z^{f}_{k}=\mathfrak{E}$, and $q_{k-1}:=\operatorname{\mathsf{E}}[x_{k-1}|\mathcal{I}^{d}_{k-1},\delta_{k-1}=1]-\operatorname{\mathsf{E}}[x_{k-1}|\mathcal{I}^{d}_{k-1}]$ when $z^{f}_{k}=\mathfrak{F}$. Then, using (25) and the definition of $p_{k-1}$, when $z^{f}_{k}=\mathfrak{E}$, i.e., when $\delta_{k-1}=0$, we get
$\operatorname{\mathsf{E}}\Big{[}x_{k}\big{|}\mathcal{I}^{d}_{k}\Big{]}=A_{k-1}\operatorname{\mathsf{E}}\Big{[}x_{k-1}\big{|}\mathcal{I}^{d}_{k-1}\Big{]}+A_{k-1}p_{k-1}$ (27)
for $k\in\mathbb{N}_{[1,N]}$, where we used the fact that $\{\mathcal{I}^{d}_{k-1},z^{f}_{k}=\mathfrak{E}\}$ is a sufficient statistic of $\mathcal{I}^{d}_{k}$. Moreover, using (25) and the definition of $q_{k-1}$, when $z^{f}_{k}=\mathfrak{F}$, i.e., when $\delta_{k-1}=1\wedge\gamma_{k-1}=0$, we get
$\operatorname{\mathsf{E}}\Big{[}x_{k}\big{|}\mathcal{I}^{d}_{k}\Big{]}=A_{k-1}\operatorname{\mathsf{E}}\Big{[}x_{k-1}\big{|}\mathcal{I}^{d}_{k-1}\Big{]}+A_{k-1}q_{k-1}$ (28)
for $k\in\mathbb{N}_{[1,N]}$, where we used the facts that $\{\mathcal{I}^{d}_{k-1},z^{f}_{k}=\mathfrak{G}\}$ is a sufficient statistic of $\mathcal{I}^{d}_{k}$ and that $x_{k-1}$ is independent of $\gamma_{k-1}$. Now, define $\imath_{k-1}:=A_{k-1}p_{k-1}$ and $\jmath_{k-1}:=A_{k-1}q_{k-1}$ as the signaling residuals. We can obtain (3) by combining (26), (27), and (28). Finally, the initial condition is $\operatorname{\mathsf{E}}[x_{0}]=m_{0}$, as no information is available at time $k=0$. ∎
The next two lemmas provide certain properties pertaining to the estimation mismatch and the signaling residuals $\imath_{k}$ and $\jmath_{k}$.
Lemma 4
The estimation mismatch satisfies the recursive equation
$\tilde{e}_{k}$ $=(1-\delta_{k-1}\gamma_{k-1})A_{k-1}\tilde{e}_{k-1}+K_{k}\nu_{k}+\mu_{k-1}$ (29)
for $k\in\mathbb{N}_{[1,N]}$ with initial condition $\tilde{e}_{0}=K_{0}\nu_{0}$, where $\mu_{k-1}=-(1-\delta_{k-1})\imath_{k-1}-\delta_{k-1}(1-\gamma_{k-1})\jmath_{k-1}$. Besides, the conditional mean and the conditional covariance of the estimation mismatch at the encoder satisfy the recursive equations
$\breve{e}_{k}$ $=(1-\bar{\theta}_{k-1})A_{k-1}\breve{e}_{k-1}+K_{k}\nu_{k}+\varsigma_{k-1}$ (30)
$R_{k}$ $=(1-\bar{\theta}_{k-1})A_{k-1}R_{k-1}A_{k-1}^{T}$
$\quad+\bar{\bar{\theta}}_{k-1}A_{k-1}\breve{e}_{k-1}\breve{e}_{k-1}^{T}A_{k-1}^{T}+\Xi_{k-1}$ (31)
for $k\in\mathbb{N}_{[1,N]}$ with initial condition $\breve{e}_{0}=K_{0}\nu_{0}$ and $R_{0}=0$, where $\bar{\theta}_{k-1}=\theta_{k-1}\delta_{k-1}\gamma_{k-1}+\lambda^{c}(1-\theta_{k-1})\delta_{k-1}$, $\bar{\bar{\theta}}_{k-1}=(1-\theta_{k-1})(\lambda^{c}\delta_{k-1}-{\lambda^{c}}^{2}\delta_{k-1})$, $\varsigma_{k-1}=\operatorname{\mathsf{E}}[\mu_{k-1}|\mathcal{I}^{e}_{k-1},\delta_{k-1},\theta_{k-1}]$, and $\Xi_{k-1}=\operatorname{\mathsf{cov}}[\mu_{k-1}|\mathcal{I}^{e}_{k-1},\delta_{k-1},\theta_{k-1}]+2\operatorname{\mathsf{cov}}[\mu_{k-1},(1-\delta_{k-1}\gamma_{k-1})A_{k-1}\tilde{e}_{k-1}|\mathcal{I}^{e}_{k-1},\delta_{k-1},\theta_{k-1}]$.
Proof:
Observe that we can obtain (29) by plugging (21) into (20), and then subtracting (3) from the result. Suppose for now that $\mu_{k-1}=0$. When feedback information is successfully delivered in the backward channel at time $k$, we have $z^{b}_{k}=\gamma_{k-1}$. Taking the conditional expectation and the conditional covariance of (29) given $\mathcal{I}^{e}_{k}$, when $z^{b}_{k}=\gamma_{k-1}$, i.e., when $\delta_{k-1}=1\wedge\theta_{k-1}=1$, we get
$\operatorname{\mathsf{E}}\Big{[}\tilde{e}_{k}\Big{|}\mathcal{I}^{e}_{k}\Big{]}$ $=(1-\sigma_{k-1})A_{k-1}\operatorname{\mathsf{E}}\Big{[}\tilde{e}_{k-1}\big{|}\mathcal{I}^{e}_{k-1}\Big{]}+K_{k}\nu_{k}$ (32)
$\operatorname{\mathsf{cov}}\Big{[}\tilde{e}_{k}\Big{|}\mathcal{I}^{e}_{k}\Big{]}$ $=(1-\sigma_{k-1})A_{k-1}\operatorname{\mathsf{cov}}\Big{[}\tilde{e}_{k-1}\big{|}\mathcal{I}^{e}_{k-1}\Big{]}A_{k-1}^{T}$ (33)
where $\sigma_{k-1}=\delta_{k-1}\gamma_{k-1}$, and we used the facts that $\{\mathcal{I}^{e}_{k-1},y_{k},z^{b}_{k}=\gamma_{k-1}\}$ is a sufficient statistic of $\mathcal{I}^{e}_{k}$, that $\tilde{e}_{k-1}$ is independent of $y_{k}$ and $z^{b}_{k}$ given $\mathcal{I}^{e}_{k-1}$, and that $\nu_{k}$ is known given $\mathcal{I}^{e}_{k}$. However, when no feedback information is delivered in the backward channel at time $k$, we have $z^{b}_{k}=\mathfrak{G}$ or $z^{b}_{k}=\mathfrak{H}$. Taking the conditional expectation and the conditional covariance of (29) given $\mathcal{I}^{e}_{k}$, when $z^{b}_{k}=\mathfrak{G}$ or $z^{b}_{k}=\mathfrak{H}$, i.e., when $\delta_{k-1}=0$ or $\delta_{k-1}=1\wedge\theta_{k-1}=0$, respectively, we get
$\operatorname{\mathsf{E}}\Big{[}\tilde{e}_{k}\big{|}\mathcal{I}^{e}_{k}\Big{]}$ $=(1-\lambda^{c}\delta_{k-1})A_{k-1}\operatorname{\mathsf{E}}\Big{[}\tilde{e}_{k-1}\big{|}\mathcal{I}^{e}_{k-1}\Big{]}+K_{k}\nu_{k}$ (34)
$\operatorname{\mathsf{cov}}\Big{[}\tilde{e}_{k}\big{|}\mathcal{I}^{e}_{k}\Big{]}$ $=(1-\lambda^{c}\delta_{k-1})A_{k-1}\operatorname{\mathsf{cov}}\Big{[}\tilde{e}_{k-1}\big{|}\mathcal{I}^{e}_{k-1}\Big{]}A_{k-1}^{T}$
$\quad+(\lambda^{c}\delta_{k-1}-{\lambda^{c}}^{2}\delta_{k-1})A_{k-1}$
×$\quad\times\operatorname{\mathsf{E}}\Big{[}\tilde{e}_{k-1}\big{|}\mathcal{I}^{e}_{k-1}\Big{]}\operatorname{\mathsf{E}}\Big{[}\tilde{e}_{k-1}\Big{|}\mathcal{I}^{e}_{k-1}\Big{]}^{T}A_{k-1}^{T}$ (35)
where we used the facts that $\{\mathcal{I}^{e}_{k-1},y_{k},z^{b}_{k}=\mathfrak{G}\}$ or $\{\mathcal{I}^{e}_{k-1},y_{k},z^{b}_{k}=\mathfrak{H}\}$ is a sufficient statistic of $\mathcal{I}^{e}_{k}$, that $\gamma_{k-1}$ is independent of $\tilde{e}_{k-1}$ and $\mathcal{I}^{e}_{k}$, that $\tilde{e}_{k-1}$ is independent of $y_{k}$ and $z^{b}_{k}$ given $\mathcal{I}^{e}_{k-1}$, and that $\nu_{k}$ is known given $\mathcal{I}^{e}_{k}$, and the properties of covariances. Now, if $\mu_{k-1}\neq 0$, it is easy to see that the contributions of $\mu_{k-1}$ in $\operatorname{\mathsf{E}}[\tilde{e}_{k}|\mathcal{I}^{e}_{k}]$ and $\operatorname{\mathsf{cov}}[\tilde{e}_{k}|\mathcal{I}^{e}_{k}]$ are captured by $\varsigma_{k-1}=\operatorname{\mathsf{E}}[\mu_{k-1}|\mathcal{I}^{e}_{k-1},\delta_{k-1},\theta_{k-1}]$ and $\Xi_{k-1}=\operatorname{\mathsf{cov}}[\mu_{k-1}|\mathcal{I}^{e}_{k-1},\delta_{k-1},\theta_{k-1}]+2\operatorname{\mathsf{cov}}[\mu_{k-1},(1-\delta_{k-1}\gamma_{k-1})A_{k-1}\tilde{e}_{k-1}|\mathcal{I}^{e}_{k-1},\delta_{k-1},\theta_{k-1}]$, respectively. As a result, following the definitions of $\breve{e}_{k}$ and $R_{k}$, we can obtain (30) by combining $\varsigma_{k-1}$ with (32) and (34), and can obtain (31) by combining $\Xi_{k-1}$ with (33) and (35). ∎
Lemma 5
Let $\operatorname{\mathsf{p}}(\delta_{k}=0|\boldsymbol{\nu}_{0:k},\boldsymbol{\delta}_{0:k-1})$ be a symmetric function of $\boldsymbol{\nu}_{0:k}$ for $k\in\mathbb{N}_{[0,N]}$. Then, $\imath_{k}$, $\jmath_{k}$, $\mu_{k}$, $\varsigma_{k}$, and $\Xi_{k}$ in Lemmas 3 and 4, are equal to zero for $k\in\mathbb{N}_{[0,N-1]}$.
Proof:
Note that, by definition, $\mu_{k}$, $\varsigma_{k}$, and $\Xi_{k}$ are zero if $\imath_{k}$ and $\jmath_{k}$ are zero. Thus, we only need to prove the claim for $\imath_{k}$ and $\jmath_{k}$ for $k\in\mathbb{N}_{[0,N-1]}$. Observe that the initial condition $\tilde{e}_{0}$ is independent of the scheduling policy. We assume that $\imath_{t}=\jmath_{t}=0$ for all $t\in\mathbb{N}_{[0,k-1]}$, and will show that $\imath_{k}=\jmath_{k}=0$. We can write
$\operatorname{\mathsf{p}}\Big{(}\boldsymbol{\delta}_{0:t}\big{|}\boldsymbol{\nu}_{0:t}\Big{)}=\prod_{t^{\prime}=0}^{t}\operatorname{\mathsf{p}}\Big{(}\delta_{t^{\prime}}\big{|}\boldsymbol{\nu}_{0:t^{\prime}},\boldsymbol{\delta}_{0:t^{\prime}-1}\Big{)}$
where we used the fact that $\delta_{t^{\prime}}$ is independent of $\boldsymbol{\nu}_{t^{\prime}+1:t}$ given $\boldsymbol{\nu}_{0:t^{\prime}}$ and $\boldsymbol{\delta}_{0:t^{\prime}-1}$. Therefore, by the hypothesis, we can deduce that $\operatorname{\mathsf{p}}(\boldsymbol{\delta}_{0:t}|\boldsymbol{\nu}_{0:t})$ is symmetric with respect to $\boldsymbol{\nu}_{0:t}$. It follows that
$\operatorname{\mathsf{p}}\Big{(}\boldsymbol{\nu}_{0:t}\big{|}\boldsymbol{\delta}_{0:t}\Big{)}\propto\operatorname{\mathsf{p}}\Big{(}\boldsymbol{\delta}_{0:t}\big{|}\boldsymbol{\nu}_{0:t}\Big{)}\operatorname{\mathsf{p}}\Big{(}\boldsymbol{\nu}_{0:t}\Big{)}.$
Since $\operatorname{\mathsf{p}}(\boldsymbol{\nu}_{0:t})$ is a symmetric distribution, we deduce that $\operatorname{\mathsf{p}}(\boldsymbol{\nu}_{0:t}|\boldsymbol{\delta}_{0:t})$ is also symmetric with respect to $\boldsymbol{\nu}_{0:t}$. For any $s\in\mathbb{N}_{[0,t]}$, by marginalization and the fact that $\boldsymbol{\nu}_{s:t}$ is independent of $\boldsymbol{\delta}_{0:s-1}$, we can find that $\operatorname{\mathsf{p}}(\boldsymbol{\nu}_{s:t}|\boldsymbol{\delta}_{s:t})$ is symmetric with respect to $\boldsymbol{\nu}_{s:t}$. Furthermore, for any $k\in\mathbb{N}_{[t+1,N]}$, we have
$\operatorname{\mathsf{p}}\Big{(}\boldsymbol{\nu}_{s:k}\big{|}\boldsymbol{\delta}_{s:t}\Big{)}=\operatorname{\mathsf{p}}\Big{(}\boldsymbol{\nu}_{s:t}\big{|}\boldsymbol{\delta}_{s:t}\Big{)}\operatorname{\mathsf{p}}\Big{(}\boldsymbol{\nu}_{t+1:k}\Big{)}.$
Therefore, we can see that $\operatorname{\mathsf{p}}(\boldsymbol{\nu}_{s:k}|\boldsymbol{\delta}_{s:t})$ is also symmetric with respect to $\boldsymbol{\nu}_{s:k}$.
Note that $z^{f}_{k+1}=\mathfrak{E}$ and $z^{f}_{k+1}=\mathfrak{F}$ are equivalent to $\delta_{k}=0$ and $\{\delta_{k}=1,\gamma_{k}=0\}$, respectively. Accordingly, when $z^{f}_{k+1}=\mathfrak{E}$ or $z^{f}_{k+1}=\mathfrak{F}$, we can write
$\operatorname{\mathsf{E}}\bigg{[}\hat{e}_{k}\Big{|}\mathcal{I}^{d}_{k},z^{f}_{k+1}\bigg{]}$
$\qquad=\operatorname{\mathsf{E}}\bigg{[}\operatorname{\mathsf{E}}\Big{[}\hat{e}_{k}\big{|}\mathcal{I}^{e}_{k},z^{f}_{k+1}\Big{]}\Big{|}\mathcal{I}^{d}_{k},z_{k+1}\bigg{]}$
$\qquad=\operatorname{\mathsf{E}}\bigg{[}\operatorname{\mathsf{E}}\Big{[}\hat{e}_{k}\big{|}\mathcal{I}^{e}_{k}\Big{]}\Big{|}\mathcal{I}^{d}_{k},z_{k+1}\bigg{]}$
$\qquad=\operatorname{\mathsf{E}}\bigg{[}\tilde{e}_{k}\Big{|}\mathcal{I}^{d}_{k},z^{f}_{k+1}\bigg{]}$
where in the first equality we used the tower property of conditional expectations, and in the second equality the facts that $\hat{e}_{k}$ is independent of $\gamma_{k}$, and that $\delta_{k}\in\mathcal{I}^{e}_{k}$. In addition, from (29), when $\imath_{t}=\jmath_{t}=0$ for all $t\in\mathbb{N}_{[d-1,k-1]}$, we can write
$\tilde{e}_{k}$ $=\sum_{t=1}^{\eta_{k}+1}\bigg{(}\prod_{t^{\prime}=1}^{t-1}A_{k-t^{\prime}}\bigg{)}K_{k-t+1}\nu_{k-t+1}$
$=D_{k}\boldsymbol{\nu}_{k-\eta_{k}:k}$ (36)
for $k\in\mathbb{N}_{[0,N]}$, where $\eta_{k}\in\mathbb{N}_{[0,k]}$ denotes the time elapsed since the last successful delivery when we are at time $k$ with the convention $\eta_{k}=k$ if no successful delivery has occurred, and $D_{k}$ is a matrix of proper dimension.
Clearly, at time $k$, the decoder knows exactly when the last successful delivery occurred, i.e., it knows $\eta_{k}$. Moreover, from the definition of $\eta_{k}$, when we are at time $k$, no successful delivery occurred from $k-\eta_{k}+1$ to $k$.
Let us define $a_{t},b_{t}\in\{0,1\}$ such that $a_{t}b_{t}=0$ for $t\in\mathbb{N}_{[0,N]}$. In the case of the channel with packet-loss detection, by (Proof:) and the definitions of $\imath_{k}$ and $\jmath_{k}$, we obtain
$\imath_{k}$ $=A_{k}D_{k}\operatorname{\mathsf{E}}\bigg{[}\boldsymbol{\nu}_{k-\eta_{k}:k}\Big{|}\mathcal{I}^{d}_{k},z^{f}_{k+1}=\mathfrak{E}\bigg{]}$
$=A_{k}D_{k}\operatorname{\mathsf{E}}\bigg{[}\boldsymbol{\nu}_{k-\eta_{k}:k}\Big{|}\mathcal{I}^{1}_{k+1}\bigg{]}$
$\jmath_{k}$ $=A_{k}D_{k}\operatorname{\mathsf{E}}\bigg{[}\boldsymbol{\nu}_{k-\eta_{k}:k}\Big{|}\mathcal{I}^{d}_{k},z^{f}_{k+1}=\mathfrak{F}\bigg{]}$
$=A_{k}D_{k}\operatorname{\mathsf{E}}\bigg{[}\boldsymbol{\nu}_{k-\eta_{k}:k}\Big{|}\mathcal{I}^{2}_{k+1}\bigg{]}$
where $\mathcal{I}^{1}_{k+1}=\{\mathcal{I}^{d}_{k-\eta_{k}},\boldsymbol{\delta}_{k-\eta_{k}:k-1}=\boldsymbol{a}_{k-\eta_{k}:k-1},\delta_{k}=0,\boldsymbol{\gamma}_{k-\eta_{k}:k-1}=\boldsymbol{b}_{k-\eta_{k}:k-1}\}$ and $\mathcal{I}^{2}_{k+1}=\{\mathcal{I}^{d}_{k-\eta_{k}},\boldsymbol{\delta}_{k-\eta_{k}:k-1}=\boldsymbol{a}_{k-\eta_{k}:k-1},\delta_{k}=1,\boldsymbol{\gamma}_{k-\eta_{k}:k-1}=\boldsymbol{b}_{k-\eta_{k}:k-1}\}$, and we used the fact that $\boldsymbol{\nu}_{k-\eta_{k}:k}$ is independent of $\gamma_{k}$. Equivalently, we can write
$\imath_{k}$ $=A_{k}D_{k}\operatorname{\mathsf{E}}\bigg{[}\boldsymbol{\nu}_{k-\eta_{k}:k}\Big{|}\bar{\mathcal{I}}^{1}_{k+1}\bigg{]}$
$\jmath_{k}$ $=A_{k}D_{k}\operatorname{\mathsf{E}}\bigg{[}\boldsymbol{\nu}_{k-\eta_{k}:k}\Big{|}\bar{\mathcal{I}}^{2}_{k+1}\bigg{]}$
where ¯$\bar{\mathcal{I}}^{1}_{k+1}=\{\boldsymbol{\delta}_{k-\eta_{k}:k-1}=\boldsymbol{a}_{k-\eta_{k}:k-1},\delta_{k}=0\}$ and ¯$\bar{\mathcal{I}}^{2}_{k+1}=\{\boldsymbol{\delta}_{k-\eta_{k}:k-1}=\boldsymbol{a}_{k-\eta_{k}:k-1},\delta_{k}=1\}$, and we used the fact that $\boldsymbol{\nu}_{k-\eta_{k}:k}$ is independent of $\mathcal{I}^{d}_{k-\eta_{k}}$ and $\boldsymbol{\gamma}_{k-\eta_{k}:k-1}$. We have already shown that $\operatorname{\mathsf{p}}(\boldsymbol{\nu}_{k-\eta_{k}:k}|\boldsymbol{\delta}_{k-\eta_{k}:k})$ is symmetric with respect to $\boldsymbol{\nu}_{k-\eta_{k}-d+1:k}$. This implies that $\imath_{k}=0$ and $\jmath_{k}=0$. ∎
The next two technical lemmas are related to symmetric decreasing rearrangements of non-negative functions. For the proofs of these lemmas, see, e.g., [58] and [59].
Lemma 6 (Hardy-Littlewood inequality)
Let $f$ and $g$ be non-negative functions defined on $\mathbb{R}^{n}$ that vanish at infinity. Then,
[…]
Lemma 7
Let $\mathcal{B}(r)\subseteq\mathbb{R}^{n}$ be a ball of radius $r$ centered at the origin, and $f$ and $g$ be non-negative functions defined on $\mathbb{R}^{n}$ that vanish at infinity and satisfy
[…]
for all $r\geq 0$. Then,
[…]
for all $r\geq 0$ and any non-negative non-increasing function $h$.
The next lemma provides an equivalent loss function.
Lemma 8
Let $\tilde{e}_{0}$ satisfies $K_{0}\nu_{0}$ and $\delta$ be constructed based on $\operatorname{\mathsf{E}}[x_{k}|\mathcal{I}^{d}_{k}]$. Optimizing the loss function $\Phi(\epsilon,\delta)$ is equivalent to optimizing the following loss function $\Psi^{N}_{\epsilon}(\tilde{e}_{0})$ over $\epsilon\in\mathcal{E}$:
$\Psi^{N}_{\epsilon}(\tilde{e}_{0})=\sum_{k=0}^{N}\Big{\{}\alpha_{k}\operatorname{\mathsf{p}}_{\epsilon}\Big{(}\boldsymbol{\sigma}_{0:k-1}=0\Big{)}$
×$\times\operatorname{\mathsf{E}}_{\epsilon}\Big{[}\delta_{k}\big{|}\boldsymbol{\sigma}_{0:k-1}=0\Big{]}$
$+\operatorname{\mathsf{p}}_{\epsilon}\Big{(}\boldsymbol{\sigma}_{0:k-1}=0\Big{)}\operatorname{\mathsf{E}}_{\epsilon}\Big{[}\tilde{e}_{k}^{T}\tilde{e}_{k}\big{|}\boldsymbol{\sigma}_{0:k-1}=0\Big{]}$
$+\operatorname{\mathsf{p}}_{\epsilon}\Big{(}\boldsymbol{\sigma}_{0:k-1}=0,\sigma_{k}=1\Big{)}$
×$\times\operatorname{\mathsf{E}}_{\epsilon}\Big{[}\Psi^{k+1,N}_{\epsilon}(\tilde{e}_{k+1})\big{|}\boldsymbol{\sigma}_{0:k-1}=0,\sigma_{k}=1\Big{]}\Big{\}}$ (40)
where $\sigma_{k}=\delta_{k}\gamma_{k}$ and
$\Psi^{k,N}_{\epsilon}(\tilde{e}_{k})=\sum_{t=k}^{N}\operatorname{\mathsf{E}}\Big{[}\alpha_{t}\delta_{t}+\tilde{e}_{t}^{T}\tilde{e}_{t}\Big{]}$
for $k\in\mathbb{N}_{[1,N]}$ when $\tilde{e}_{k}$ satisfies $K_{k}\nu_{k}$.
Proof:
From the definition of the loss function $\Phi(\epsilon,\delta)$, we can write
$\Phi(\epsilon,\delta)$ $=\sum_{k=0}^{N}\operatorname{\mathsf{E}}\bigg{[}\alpha_{k}\delta_{k}+\hat{e}_{k}^{T}\hat{e}_{k}\bigg{]}$
$=\sum_{k=0}^{N}\operatorname{\mathsf{E}}\bigg{[}\operatorname{\mathsf{E}}\Big{[}\alpha_{k}\delta_{k}+\hat{e}_{k}^{T}\hat{e}_{k}\big{|}\mathcal{I}^{e}_{k}\Big{]}\bigg{]}$
$=\sum_{k=0}^{N}\operatorname{\mathsf{E}}\bigg{[}\alpha_{k}\delta_{k}+\tilde{e}_{k}^{T}\tilde{e}_{k}+\operatorname{tr}Q_{k}\bigg{]}$ (41)
where in the second equality we used the tower property of conditional expectations. Observe that, by Lemma 4, we have $\tilde{e}_{k+1}=K_{k+1}\nu_{k+1}$ only when $\sigma_{k}=1$. We can show that, associated with the optimal scheduling policy, the following relation is satisfied:
$\operatorname{\mathsf{E}}\bigg{[}\alpha_{t}\delta_{t}+\tilde{e}_{t}^{T}\tilde{e}_{t}\Big{|}\sigma_{k}=1\bigg{]}$
$\quad=\operatorname{\mathsf{E}}\bigg{[}\alpha_{t}\delta_{t}+\tilde{e}_{t}^{T}\tilde{e}_{t}\Big{|}\boldsymbol{\sigma}_{0:k-d},\sigma_{k}=1\bigg{]}$ (42)
for any $t\in\mathbb{N}_{[k+1,N]}$, where we used the fact that, without loss of optimality, $\boldsymbol{\delta}_{k+1:N}$ and $\tilde{\boldsymbol{e}}_{k+1:N}$ are independent of $\boldsymbol{\sigma}_{0:k-1}$ given $\sigma_{k}=1$, i.e., when $\tilde{e}_{k+1}=K_{k+1}\nu_{k+1}$.
Now, define the loss function $\Psi^{N}_{\epsilon}(\tilde{e}_{0})$ as
$\Psi^{N}_{\epsilon}(\tilde{e}_{0}):=\sum_{k=0}^{N}\operatorname{\mathsf{E}}\bigg{[}\alpha_{k}\delta_{k}+\tilde{e}_{k}^{T}\tilde{e}_{k}\bigg{]}.$ (43)
Following the facts that $\tilde{e}_{0}$ and $\operatorname{tr}Q_{t}$ for $t\in\mathbb{N}_{[0,N]}$ are independent of the scheduling policy, to optimize $\Phi(\epsilon,\delta)$ in (Proof:) over $\epsilon\in\mathcal{E}$, it suffices to optimize $\Psi^{N}_{\epsilon}(\tilde{e}_{0})$ in (43) over $\epsilon\in\mathcal{E}$. In addition, observe that from the law of total probability, we have
$\operatorname{\mathsf{p}}_{\epsilon}\Big{(}\sigma_{0}=1\Big{)}+\operatorname{\mathsf{p}}_{\epsilon}\Big{(}\boldsymbol{\sigma}_{0:t}=0\Big{)}$
$\qquad\qquad+\sum_{t^{\prime}=1}^{t}\operatorname{\mathsf{p}}_{\epsilon}\Big{(}\boldsymbol{\sigma}_{0:t^{\prime}-1}=0,\sigma_{t^{\prime}}=1\Big{)}=1$ (44)
for any $t\in\mathbb{N}_{[0,N]}$. Applying the law of total expectation for the terms $\operatorname{\mathsf{E}}[\alpha_{k}\delta_{k}]$ and $\operatorname{\mathsf{E}}[\tilde{e}_{k}^{T}\tilde{e}_{k}]$ in $\Psi^{N}_{\epsilon}(\tilde{e}_{0})$ on a partition provided by the identity (Proof:) for $t=k-1$, repeating this procedure for $k\in\mathbb{N}_{[1,N]}$, and using the definition of $\Psi^{k+1,N}_{\epsilon}(\tilde{e}_{k+1})$ along with the relation in (Proof:), we obtain (LABEL:eq:lemma:Omega). This completes the proof.
In light of Lemma 1 and from the definition of the loss function $\Phi(\epsilon,\delta)$, we can write
$\Phi(\epsilon,\delta)$ $=\sum_{k=0}^{N}\operatorname{\mathsf{E}}\Big{[}\alpha_{k}\delta_{k}+\hat{e}_{k}^{T}\hat{e}_{k}\Big{]}$
$=\sum_{k=0}^{N}\operatorname{\mathsf{E}}\Big{[}\alpha_{k}\delta_{k}+\tilde{e}_{k}^{T}\tilde{e}_{k}+\operatorname{tr}Q_{k}\Big{]}$
where $\delta$ is chosen based on $\hat{x}_{k}=\operatorname{\mathsf{E}}[x_{k}|\mathcal{I}^{d}_{k}]$, and in the second equality we used the tower property of conditional expectations. Define the loss function $\Psi^{N}_{\epsilon}(\tilde{e}_{0})$ as
$\Psi^{N}_{\epsilon}(\tilde{e}_{0}):=\sum_{k=0}^{N}\operatorname{\mathsf{E}}\Big{[}\alpha_{k}\delta_{k}+\tilde{e}_{k}^{T}\tilde{e}_{k}\Big{]}.$
Following the fact that $\operatorname{tr}Q_{k}$ is independent of $\epsilon$ and $\delta$, to optimize $\Phi(\epsilon,\delta)$ over $\epsilon\in\mathcal{E}$, it suffices to optimize over $\epsilon\in\mathcal{E}$. In addition, observe that from the law of total probability, we have
$\operatorname{\mathsf{p}}_{\epsilon}\Big{(}\sigma_{0}=1\Big{)}+\operatorname{\mathsf{p}}_{\epsilon}\Big{(}\boldsymbol{\sigma}_{0:t}=0\Big{)}$
$\qquad\qquad+\sum_{t^{\prime}=1}^{t}\operatorname{\mathsf{p}}_{\epsilon}\Big{(}\boldsymbol{\sigma}_{0:t^{\prime}-1}=0,\sigma_{t^{\prime}}=1\Big{)}=1$ (45)
for any $t\in\mathbb{N}_{[0,N]}$. Applying the law of total expectation for the terms $\operatorname{\mathsf{E}}[\alpha_{k}\delta_{k}]$ and $\operatorname{\mathsf{E}}[\tilde{e}_{k}^{T}\tilde{e}_{k}]$ in $\Psi^{N}_{\epsilon}(\tilde{e}_{0})$ on a partition provided by the identity (Proof:) for $t=k-1$, repeating this procedure for $k\in\mathbb{N}_{[1,N]}$, and using the definition of $\Psi^{k,N}_{\epsilon}(\tilde{e}_{k})$, we can obtain (8). ∎
We are now in a position to provide the proof of Theorem 1.
Proof:
Suppose that $(\epsilon^{o},\delta^{o})$ is a coding policy profile that belongs to the set of globally optimal solutions, which obviously cannot be empty. We prove global optimality of the policy profile $(\epsilon^{\star},\delta^{\star})$ in the claim by showing that $\Phi(\epsilon^{\star},\delta^{\star})$ cannot be greater than $\Phi^{o}:=\Phi(\epsilon^{o},\delta^{o})$. Without loss of generality, we assume that $m_{0}=0$. Similar arguments can be made for $m_{0}\neq 0$ following a coordinate transformation. The proof consists of three steps.
Step 1. We say a scheduling policy is innovation-based if, at each time $k$, it depends on $\boldsymbol{\nu}_{0:k}$ instead of $\boldsymbol{y}_{0:k}$ and $\boldsymbol{z}^{b}_{0:k}$, i.e., if it can be expressed in the form $\operatorname{\mathsf{p}}(\delta_{k}|\boldsymbol{\nu}_{0:k},\boldsymbol{\delta}_{0:k-1},\boldsymbol{\theta}_{0:k-1})$. In the first step of our proof, we show that $\Phi(\epsilon^{n},\delta^{o})=\Phi(\epsilon^{o},\delta^{o})$, where $\epsilon^{n}$ is an innovation-based scheduling policy. From the definition of the innovation, (20), and (21), we can write $\boldsymbol{y}_{0:k}=\boldsymbol{\nu}_{0:k}+G_{k}\check{\boldsymbol{x}}_{0:k-1}$ and $\check{\boldsymbol{x}}_{0:k}=H_{k}\boldsymbol{\nu}_{0:k}$, where $G_{k}$ and $H_{k}$ are matrices of proper dimensions. Putting these equations together, we find $\boldsymbol{y}_{0:k}=\boldsymbol{\nu}_{0:k}+G_{k}H_{k-1}\boldsymbol{\nu}_{0:k-1}$. Therefore, $\operatorname{\mathsf{p}}_{\epsilon^{o}}(\delta_{k}|\boldsymbol{y}_{0:k},\boldsymbol{z}^{b}_{0:k},\boldsymbol{\delta}_{0:k-1},\boldsymbol{\theta}_{0:k-1})$ can equivalently be written as $\operatorname{\mathsf{p}}_{\epsilon^{n}}(\delta_{k}|\boldsymbol{\nu}_{0:k},\boldsymbol{z}^{b}_{0:k},\boldsymbol{\delta}_{0:k-1},\boldsymbol{\theta}_{0:k-1})$. This establishes that $\Phi(\epsilon^{n},\delta^{o})=\Phi(\epsilon^{o},\delta^{o})$. As our subsequent analysis is valid for any values of $\boldsymbol{z}^{b}_{0:k}$ and $\boldsymbol{\theta}_{0:k-1}$, for brevity, hereafter we omit the dependency of $\epsilon^{n}$ on these variables.
Step 2. Let $\mathcal{B}(r)$ be a ball of radius $r$ centered at the origin and of proper dimension. Define $\varpi_{k}:=T_{k}\boldsymbol{\nu}_{0:k}\in\mathcal{N}$ for a specific matrix $T_{k}$, and $\sigma_{k}:=\delta_{k}\gamma_{k}$, $l_{\epsilon^{s}}(\varpi_{k}):=(\lambda+\lambda^{c}\operatorname{\mathsf{p}}_{\epsilon^{s}}(\delta_{k}=0|\varpi_{k},\boldsymbol{\sigma}_{0:k-1}=0))\operatorname{\mathsf{p}}_{\epsilon^{s}}(\varpi_{k}|\boldsymbol{\sigma}_{0:k-1}=0)$, and $l_{\epsilon^{n}}(\varpi_{k}):=(\lambda+\lambda^{c}\operatorname{\mathsf{p}}_{\epsilon^{n}}(\delta_{k}=0|\varpi_{k},\boldsymbol{\sigma}_{0:k-1}=0))\operatorname{\mathsf{p}}_{\epsilon^{n}}(\varpi_{k}|\boldsymbol{\sigma}_{0:k-1}=0)$. In the second step of our proof, we show that $\Phi(\epsilon^{s},\delta^{o})\leq\Phi(\epsilon^{n},\delta^{o})$, where $\epsilon^{s}$ is a special form of $\epsilon^{n}$ that is symmetric with respect to $\boldsymbol{\nu}_{0:k}$ at each time $k$ and such that the following conditions are satisfied:
[…]
for all $r\geq 0$ with $l_{\epsilon^{s}}(\varpi_{k})$ as a radially symmetric function of $\varpi_{k}$. We know that $\sigma_{k}=0$ either when $\delta_{k}=1\;\wedge\;\gamma_{k}=0$ or when $\delta_{k}=0$ regardless of $\gamma_{k}$. Accordingly, we can write
$\operatorname{\mathsf{p}}_{\epsilon^{n}}\Big{(}\sigma_{k}=0\big{|}\varpi_{k},\boldsymbol{\sigma}_{0:k-1}=0\Big{)}$
$\ =\operatorname{\mathsf{p}}_{\epsilon^{n}}\Big{(}\delta_{k}=1\big{|}\varpi_{k},\boldsymbol{\sigma}_{0:k-1}=0)$
×$\qquad\times\operatorname{\mathsf{p}}_{\epsilon^{n}}\Big{(}\delta_{k}\gamma_{k}=0\big{|}\varpi_{k},\boldsymbol{\sigma}_{0:k-1}=0,\delta_{k}=1\Big{)}$
$\qquad+\operatorname{\mathsf{p}}_{\epsilon^{n}}\Big{(}\delta_{k}=0\big{|}\varpi_{k},\boldsymbol{\sigma}_{0:k-1}=0\Big{)}$
×$\qquad\times\operatorname{\mathsf{p}}_{\epsilon^{n}}\Big{(}\delta_{k}\gamma_{k}=0\big{|}\varpi_{k},\boldsymbol{\sigma}_{0:k-1}=0,\delta_{k}=0\Big{)}$
$\ =\lambda\operatorname{\mathsf{p}}_{\epsilon^{n}}\Big{(}\delta_{k}=1\big{|}\varpi_{k},\boldsymbol{\sigma}_{0:k-1}=0\Big{)}$
$\qquad+\operatorname{\mathsf{p}}_{\epsilon^{n}}\Big{(}\delta_{k}=0\big{|}\varpi_{k},\boldsymbol{\sigma}_{0:k-1}=0\Big{)}$
$\ =\lambda+\lambda^{c}\operatorname{\mathsf{p}}_{\epsilon^{n}}\Big{(}\delta_{k}=0\big{|}\varpi_{k},\boldsymbol{\sigma}_{0:k-1}=0\Big{)}$
where in the first equality we used the law of total probability and in the second equality the fact that $\gamma_{k}$ is independent of $\varpi_{k}$, $\boldsymbol{\sigma}_{0:k-1}$, and $\delta_{k}$. Define $l^{\prime}_{\epsilon^{s}}(\varpi_{k})=\operatorname{\mathsf{p}}_{\epsilon^{s}}(\sigma_{k}=0|\varpi_{k},\boldsymbol{\sigma}_{0:k-1}=0)\operatorname{\mathsf{p}}_{\epsilon^{s}}(\varpi_{k}|\boldsymbol{\sigma}_{0:k-1}=0)$ and $l^{\prime}_{\epsilon^{s}}(\varpi_{k})=\operatorname{\mathsf{p}}_{\epsilon^{s}}(\sigma_{k}=0|\varpi_{k},\boldsymbol{\sigma}_{0:k-1}=0)\operatorname{\mathsf{p}}_{\epsilon^{s}}(\varpi_{k}|\boldsymbol{\sigma}_{0:k-1}=0)$. Therefore, the above condition can be written in the following equivalent form:
[…]
for all $r\geq 0$ with $l^{\prime}_{\epsilon^{s}}(\varpi_{k})$ as a radially symmetric function of $\varpi_{k}$. By Lemma 8, to prove that $\Phi(\epsilon^{s},\delta^{o})\leq\Phi(\epsilon^{n},\delta^{o})$, it suffices to prove that $\Psi^{M}_{\epsilon^{s}}(\tilde{e}_{0})\leq\Psi^{M}_{\epsilon^{n}}(\tilde{e}_{0})$ for any $M\in\mathbb{N}_{[0,N]}$ and for any $\tilde{e}_{0}$. Observe that $\tilde{e}_{0}=K_{0}\nu_{0}$ is the same under both $\epsilon^{n}$ and $\epsilon^{s}$, and that $\delta_{0}$ has no effects on remote state estimation when the time horizon is 00. Hence, the claim holds for the time horizon is 00. We assume that the claim also holds for all time horizons from $1$ to $M-1$, and will show that the terms in $\Psi^{M}_{\epsilon^{n}}(\tilde{e}_{0})$ are not less than those in $\Psi^{M}_{\epsilon^{s}}(\tilde{e}_{0})$.
Step 2.1 First, for the probability coefficients, we have
$\operatorname{\mathsf{p}}_{\epsilon^{n}}\Big{(}\sigma_{k-1}=0\big{|}\boldsymbol{\sigma}_{0:k-2}=0\Big{)}$
$=\int_{\mathbb{R}^{m}}\operatorname{\mathsf{p}}_{\epsilon^{n}}(\sigma_{k-1}=0|\varpi_{k-1},\boldsymbol{\sigma}_{0:k-2}=0)$
×$\qquad\qquad\qquad\times\operatorname{\mathsf{p}}_{\epsilon^{n}}(\varpi_{k-1}|\boldsymbol{\sigma}_{0:k-2}=0)d\varpi_{k-1}$
$=\int_{\mathbb{R}^{m}}\operatorname{\mathsf{p}}_{\epsilon^{s}}(\sigma_{k-1}=0|\varpi_{k-1},\boldsymbol{\sigma}_{0:k-2}=0)$
×$\qquad\qquad\qquad\times\operatorname{\mathsf{p}}_{\epsilon^{s}}(\varpi_{k-1}|\boldsymbol{\sigma}_{0:k-2}=0)d\varpi_{k-1}$
$=\operatorname{\mathsf{p}}_{\epsilon^{s}}\Big{(}\sigma_{k-1}=0\big{|}\boldsymbol{\sigma}_{0:k-2}=0\Big{)}$
[…]
Step 2.2. Moreover, for the terms involving the expected value of the transmission decision, we have
$\operatorname{\mathsf{E}}_{\epsilon^{n}}\Big{[}$ $\delta_{k}\big{|}\boldsymbol{\sigma}_{0:k-1}=0\Big{]}$
$=1-\operatorname{\mathsf{p}}_{\epsilon^{n}}(\delta_{k}=0|\boldsymbol{\sigma}_{0:k-1}=0)$
$=\frac{1}{\lambda^{c}}-\frac{1}{\lambda^{c}}\operatorname{\mathsf{p}}_{\epsilon^{n}}(\sigma_{k}=0|\boldsymbol{\sigma}_{0:k-1}=0)$
$=\frac{1}{\lambda^{c}}-\frac{1}{\lambda^{c}}\operatorname{\mathsf{p}}_{\epsilon^{s}}(\sigma_{k}=0|\boldsymbol{\sigma}_{0:k-1}=0)$
$=1-\operatorname{\mathsf{p}}_{\epsilon^{s}}(\delta_{k}=0|\boldsymbol{\sigma}_{0:k-1}=0)$
$=\operatorname{\mathsf{E}}_{\epsilon^{s}}\Big{[}\delta_{k}\big{|}\boldsymbol{\sigma}_{0:k-1}=0\Big{]}.$
Step 2.3. Observe that by Lemma 4, when $\boldsymbol{\sigma}_{0:k-1}=0$, $\tilde{e}_{t}$ satisfies the recursive equation
$\tilde{e}_{t}$ $=A_{t-1}\tilde{e}_{t-1}+K_{t}\nu_{t}-(1-\delta_{t-1})\imath_{t-1}-\delta_{t-1}\jmath_{t-1}$
for $t\in\mathbb{N}_{[1,k]}$ with initial condition $\tilde{e}_{0}=K_{0}\nu_{0}$. Accordingly, we can find a proper matrix $E_{k}$ and a proper vector $c_{k}$, both independent of $\boldsymbol{\nu}_{0:k}$, such that $\tilde{e}_{k}=E_{k}\boldsymbol{\nu}_{0:k-1}+K_{k}\nu_{k}+c_{k}$ under $\epsilon^{n}$. We know that $\epsilon^{s}$ is symmetric with respect to $\boldsymbol{\nu}_{0:k}$ at each time $k$. Therefore, by Lemma 5, we deduce that $\tilde{e}_{k}=E_{k}\boldsymbol{\nu}_{0:k-1}+K_{k}\nu_{k}$ under $\epsilon^{s}$. For the terms involving the expected value of the quadratic estimation mismatch, we can then write
$\operatorname{\mathsf{E}}_{\epsilon^{n}}\Big{[}\tilde{e}_{k}^{T}\tilde{e}_{k}\big{|}\boldsymbol{\sigma}_{0:k-1}=0\Big{]}$
$\ =\operatorname{\mathsf{E}}_{\epsilon^{n}}\Big{[}\big{(}E_{k}\boldsymbol{\nu}_{0:k-1}+K_{k}\nu_{k}+c_{k}\big{)}^{T}$
×$\qquad\times\big{(}E_{k}\boldsymbol{\nu}_{0:k-1}+K_{k}\nu_{k}+c_{k}\big{)}\big{|}\boldsymbol{\sigma}_{0:k-1}=0\Big{]}$
$\ =\operatorname{\mathsf{E}}_{\epsilon^{n}}\Big{[}\big{(}E_{k}\boldsymbol{\nu}_{0:k-1}+c_{k}\big{)}^{T}\big{(}E_{k}\boldsymbol{\nu}_{0:k-1}+c_{k}\big{)}$
$\qquad+\nu_{k}^{T}K_{k}^{T}K_{k}\nu_{k}\big{|}\boldsymbol{\sigma}_{0:k-1}=0\Big{]}$
where in the second equality we used the fact that $\nu_{k}$ has zero mean and is independent of $\boldsymbol{\nu}_{0:k-1}$ and $\boldsymbol{\sigma}_{0:k-1}$. Choose $T_{k-1}=E_{k}$, and define $f_{\epsilon^{n}}(\varpi_{k-1},\nu_{k}):=(\varpi_{k-1}+c_{k})^{T}(\varpi_{k-1}+c_{k})+\nu_{k}^{T}K_{k}^{T}K_{k}\nu_{k}$, $f_{\epsilon^{s}}(\varpi_{k-1},\nu_{k}):=\varpi_{k-1}^{T}\varpi_{k-1}+\nu_{k}^{T}K_{k}^{T}K_{k}\nu_{k}$, $g_{\epsilon^{n}}(\varpi_{k-1},\nu_{k}):=z-\min_{z}\{z,f_{\epsilon^{n}}(\varpi_{k-1},\nu_{k})\}$, and $g_{\epsilon^{s}}(\varpi_{k-1},\nu_{k}):=z-\min_{z}\{z,f_{\epsilon^{s}}(\varpi_{k-1},\nu_{k})\}$. Clearly, $g_{\epsilon^{n}}(\varpi_{k-1},\nu_{k})$ and $g_{\epsilon^{s}}(\varpi_{k-1},\nu_{k})$ both vanish at infinity for any fixed $z$. It follows that
$\operatorname{\mathsf{E}}_{\epsilon^{n}}\Big{[}\tilde{e}_{k}^{T}\tilde{e}_{k}\big{|}\boldsymbol{\sigma}_{0:k-1}=0\Big{]}$
$\qquad=\int_{\mathbb{R}^{m}}\int_{\mathbb{R}^{m}}f_{\epsilon^{n}}(\varpi_{k-1},\nu_{k})\operatorname{\mathsf{p}}_{\epsilon^{n}}(\varpi_{k-1}|\boldsymbol{\sigma}_{0:k-1}=0)$
×$\qquad\qquad\qquad\qquad\times\operatorname{\mathsf{p}}(\nu_{k})d\varpi_{k-1}d\nu_{k}.$
In addition, we can write
$\int_{\mathbb{R}^{m}}g_{\epsilon^{n}}(\varpi_{k-1},\nu_{k})\operatorname{\mathsf{p}}_{\epsilon^{s}}(\varpi_{k-1}|\boldsymbol{\sigma}_{0:k-2}=0)$
×$\qquad\times\operatorname{\mathsf{p}}_{\epsilon^{n}}(\sigma_{k-1}=0|\varpi_{k-1},\boldsymbol{\sigma}_{0:k-2}=0)d\varpi_{k-1}$
$\leq\int_{\mathbb{R}^{m}}g^{*}_{\epsilon^{n}}(\varpi_{k-1},\nu_{k})\big{(}\operatorname{\mathsf{p}}_{\epsilon^{s}}(\varpi_{k-1}|\boldsymbol{\sigma}_{0:k-2}=0)$
×$\qquad\times\operatorname{\mathsf{p}}_{\epsilon^{n}}(\sigma_{k-1}=0|\varpi_{k-1},\boldsymbol{\sigma}_{0:k-2}=0)\big{)}^{*}d\varpi_{k-1}$
$=\int_{\mathbb{R}^{m}}g_{\epsilon^{s}}(\varpi_{k-1},\nu_{k})\big{(}\operatorname{\mathsf{p}}_{\epsilon^{s}}(\varpi_{k-1}|\boldsymbol{\sigma}_{0:k-2}=0)$
×$\qquad\times\operatorname{\mathsf{p}}_{\epsilon^{n}}(\sigma_{k-1}=0|\varpi_{k-1},\boldsymbol{\sigma}_{0:k-2}=0)\big{)}^{*}d\varpi_{k-1}$
$\leq\int_{\mathbb{R}^{m}}g_{\epsilon^{s}}(\varpi_{k-1},\nu_{k})\operatorname{\mathsf{p}}_{\epsilon^{n}}(\varpi_{k-1}|\boldsymbol{\sigma}_{0:k-2}=0)$
×$\qquad\times\operatorname{\mathsf{p}}_{\epsilon^{s}}(\sigma_{k-1}=0|\varpi_{k-1},\boldsymbol{\sigma}_{0:k-2}=0)d\varpi_{k-1}$
where in the first inequality we used the Hardy-Littlewood inequality with respect to $\varpi_{k-1}$, in the equality the fact that $g^{*}_{\epsilon^{n}}(\varpi_{k-1},\nu_{k})=g_{\epsilon^{s}}(\varpi_{k-1},\nu_{k})$, and in the second inequality (47) and Lemma 7. This implies that
[…]
×$\qquad\qquad\qquad\times\operatorname{\mathsf{p}}_{\epsilon^{n}}(\varpi_{k-1}|\boldsymbol{\sigma}_{0:k-1}=0)d\varpi_{k-1}$
[…]
×$\qquad\qquad\qquad\times\operatorname{\mathsf{p}}_{\epsilon^{s}}(\varpi_{k-1}|\boldsymbol{\sigma}_{0:k-1}=0)d\varpi_{k-1}$ (48)
where we applied the Bayes’ rule on $\operatorname{\mathsf{p}}_{\epsilon^{n}}(\varpi_{k-1}|\boldsymbol{\sigma}_{0:k-1}=0)$ and used the fact that $\operatorname{\mathsf{p}}_{\epsilon^{n}}(\sigma_{k-1}=0|\boldsymbol{\sigma}_{0:k-2}=0)=\operatorname{\mathsf{p}}_{\epsilon^{s}}(\sigma_{k-1}=0|\boldsymbol{\sigma}_{0:k-2}=0)$. Then, we deduce that
$\operatorname{\mathsf{E}}_{\epsilon^{n}}\Big{[}\tilde{e}_{k}^{T}\tilde{e}_{k}\big{|}\boldsymbol{\sigma}_{0:k-1}=0\Big{]}$
$\quad=\int_{\mathbb{R}^{m}}\int_{\mathbb{R}^{m}}f_{\epsilon^{n}}(\varpi_{k-1},\nu_{k})\operatorname{\mathsf{p}}_{\epsilon^{n}}(\varpi_{k-1}|\boldsymbol{\sigma}_{0:k-1}=0)$
×$\qquad\qquad\qquad\qquad\times\operatorname{\mathsf{p}}(\nu_{k})d\varpi_{k-1}d\nu_{k}$
$\quad\geq\int_{\mathbb{R}^{m}}\int_{\mathbb{R}^{m}}f_{\epsilon^{s}}(\varpi_{k-1},\nu_{k})\operatorname{\mathsf{p}}_{\epsilon^{s}}(\varpi_{k-1}|\boldsymbol{\sigma}_{0:k-1}=0)$
×$\qquad\qquad\qquad\qquad\times\operatorname{\mathsf{p}}(\nu_{k})d\varpi_{k-1}d\nu_{k}$
$\quad=\operatorname{\mathsf{E}}_{\epsilon^{s}}\Big{[}\tilde{e}_{k}^{T}\tilde{e}_{k}\big{|}\boldsymbol{\sigma}_{0:k-1}=0\Big{]}$
where the inequality is obtained from (Proof:) after taking $z$ to infinity.
Step 2.4. In addition, for the terms involving the expected value of the cost-to-go, we find
$\operatorname{\mathsf{E}}_{\epsilon^{n}}\Big{[}\Psi^{k+1,M}_{\epsilon^{n}}(\tilde{e}_{k+1})\big{|}\boldsymbol{\sigma}_{0:k-1}=0,\sigma_{k}=1\Big{]}$
$\quad=\int_{\mathbb{R}^{m\times(k+2)}}\Psi^{k+1,M}_{\epsilon^{n}}(\tilde{e}_{k+1})$
×$\qquad\qquad\times\operatorname{\mathsf{p}}_{\epsilon^{n}}(\boldsymbol{\nu}_{0:k+1}|\boldsymbol{\sigma}_{0:k-1}=0,\sigma_{k}=1)d\boldsymbol{\nu}_{0:k+1}.$
By Lemma 4, when $\sigma_{k}=1$, we have $\tilde{e}_{k+1}=K_{k+1}\nu_{k+1}$. Hence, $\tilde{e}_{k+1}$ is the same under both $\epsilon^{n}$ and $\epsilon^{s}$. For any $M$, let $\bar{\Psi}^{M}_{\epsilon^{n}}(\tilde{e}_{0})$ represent a loss function that is structurally similar to $\Psi^{M}_{\epsilon^{n}}(\tilde{e}_{0})$ but with new values of $\boldsymbol{\alpha}_{0:M}$, say $\bar{\boldsymbol{\alpha}}_{0:M}$. Clearly, if $\Psi^{M}_{\epsilon^{n}}(\tilde{e}_{0})\geq\Psi^{M}_{\epsilon^{s}}(\tilde{e}_{0})$, then $\bar{\Psi}^{M}_{\epsilon^{n}}(\tilde{e}_{0})\geq\bar{\Psi}^{M}_{\epsilon^{s}}(\tilde{e}_{0})$. Accordingly, we can write
$\int_{\mathbb{R}^{m\times(k+2)}}\Psi^{k+1,M}_{\epsilon^{n}}(\tilde{e}_{k+1})$
×$\qquad\quad\times\operatorname{\mathsf{p}}_{\epsilon^{n}}(\boldsymbol{\nu}_{0:k+1}|\boldsymbol{\sigma}_{0:k-1}=0,\sigma_{k}=1)d\boldsymbol{\nu}_{0:k+1}$
[…]
$=\int_{\mathbb{R}^{m\times(k+2)}}\Psi^{k+1,M}_{\epsilon^{s}}(\tilde{e}_{k+1})$
×$\qquad\quad\times\operatorname{\mathsf{p}}_{\epsilon^{s}}(\boldsymbol{\nu}_{0:k+1}|\boldsymbol{\sigma}_{0:k-1}=0,\sigma_{k}=1)d\boldsymbol{\nu}_{0:k+1}$
where in the equalities we used the facts that $\Psi^{k+1,M}_{\epsilon^{n}}(\tilde{e})=\bar{\Psi}^{M-k-1}_{\epsilon^{n}}(\tilde{e})$ for any Gaussian variable $\tilde{e}$ and a suitable selection of $\bar{\boldsymbol{\alpha}}_{0:M-k-1}$, and that $\nu_{k+1}$ is independent of $\boldsymbol{\sigma}_{0:k}$, and the Fubini’s theorem; and in the inequality we used the hypothesis $\Psi^{M-k-1}_{\epsilon^{n}}(\tilde{e})\geq\Psi^{M-k-1}_{\epsilon^{s}}(\tilde{e})$ for any Gaussian variable $\tilde{e}$. Therefore,
$\operatorname{\mathsf{E}}_{\epsilon^{n}}\Big{[}\Psi^{k+1,M}_{\epsilon^{n}}(\tilde{e}_{k+1})\big{|}\boldsymbol{\sigma}_{0:k-1}=0,\sigma_{k}=1\Big{]}$
$\qquad\geq\operatorname{\mathsf{E}}_{\epsilon^{s}}\Big{[}\Psi^{k+1,M}_{\epsilon^{s}}(\tilde{e}_{k+1})\big{|}\boldsymbol{\sigma}_{0:k-1}=0,\sigma_{k}=1\Big{]}.$
This establishes that $\Psi^{M}_{\epsilon^{s}}(\tilde{e}_{0})\leq\Psi^{M}_{\epsilon^{n}}(\tilde{e}_{0})$, and verifies that $\Phi(\epsilon^{s},\delta^{o})\leq\Phi(\epsilon^{n},\delta^{o})$.
Step 3. Now, in the final step of our proof, we show that $\Phi(\epsilon^{\star},\delta^{\star})\leq\Phi(\epsilon^{s},\delta^{o})$, where $\epsilon^{\star}$ is a special form of $\epsilon^{s}$ and $\delta^{\star}$ is the same as $\delta^{o}$. Observe that, by Lemma 1 and Lemmas 3 and 5, when $\epsilon^{s}$ is used, $\delta^{o}$ must satisfy the recursive equation
$\hat{x}_{k}=A_{k-1}\hat{x}_{k-1}+\delta_{k-1}\gamma_{k-1}A_{k-1}(\check{x}_{k-1}-\hat{x}_{k-1})$ (49)
for $k\in\mathbb{N}_{[1,N]}$ with initial condition $\hat{x}_{0}=m_{0}=0$. Moreover, from the definition of the value function $V_{k}(\mathcal{I}^{e}_{k})$ in (13), we can write
$V_{k}(\mathcal{I}^{e}_{k})=\min_{\delta_{k}\in\{0,1\}}\bigg{\{}\alpha_{k}\delta_{k}+\operatorname{\mathsf{E}}[\tilde{e}_{k+1}^{T}\tilde{e}_{k+1}|\mathcal{I}^{e}_{k}]$
$\qquad\qquad\qquad\qquad+\operatorname{tr}Q_{k+1}+\operatorname{\mathsf{E}}[V_{k+1}(\mathcal{I}^{e}_{k+1})|\mathcal{I}^{e}_{k}]\bigg{\}}$ (50)
for $k\in\mathbb{N}_{[0,N-1]}$ with initial condition $V_{N}(\mathcal{I}^{e}_{N})=0$, where we used the additivity property of $V_{k}(\mathcal{I}^{e}_{k})$ and the fact that $\operatorname{\mathsf{E}}[\hat{e}_{k+1}^{T}\hat{e}_{k+1}|\mathcal{I}^{e}_{k}]=\operatorname{\mathsf{E}}[\tilde{e}_{k+1}^{T}\tilde{e}_{k+1}|\mathcal{I}^{e}_{k}]+\operatorname{tr}Q_{k+1}$ according to the tower properties of conditional expectations. The minimizing scheduling policy is then obtained as
$\delta_{k}=\mathds{1}_{\chi_{k}-\alpha_{k}\geq 0}$ (51)
where $\chi_{k}=\operatorname{\mathsf{E}}[\tilde{e}_{k+1}^{T}\tilde{e}_{k+1}+V_{k+1}(\mathcal{I}^{e}_{k+1})|\mathcal{I}^{e}_{k},\delta_{k}=0]-\operatorname{\mathsf{E}}[\tilde{e}_{k+1}^{T}\tilde{e}_{k+1}+V_{k+1}(\mathcal{I}^{e}_{k+1})|\mathcal{I}^{e}_{k},\delta_{k}=1]$. Next, we prove by induction that $V^{e}_{k}(\mathcal{I}^{e}_{k})$ is a symmetric function of $\breve{e}_{k}$ and a function $R_{k}$. Clearly, the claim holds for time $N$. We assume that the claim holds for time $k+1$ and we will show that it also holds for time $k$. Let us write $\chi_{k}=\chi^{1}_{k}+\chi^{2}_{k}$, where $\chi^{1}_{k}=\operatorname{\mathsf{E}}[\tilde{e}_{k+1}^{T}\tilde{e}_{k+1}|\mathcal{I}^{e}_{k},\delta_{k}=0]-\operatorname{\mathsf{E}}[\tilde{e}_{k+1}^{T}\tilde{e}_{k+1}|\mathcal{I}^{e}_{k},\delta_{k}=1]$ and $\chi^{2}_{k}=\operatorname{\mathsf{E}}[V_{k+1}(\breve{e}_{k+1},R_{k+1})|\mathcal{I}^{e}_{k},\delta_{k}=0]-\operatorname{\mathsf{E}}[V_{k+1}(\breve{e}_{k+1},R_{k+1})|\mathcal{I}^{e}_{k},\delta_{k}=1]$. From the properties of expectations, we have the following relation
$\operatorname{\mathsf{E}}\Big{[}\tilde{e}_{k+1}^{T}\tilde{e}_{k+1}\big{|}\mathcal{I}^{e}_{k},\delta_{k}\Big{]}$ $=\operatorname{\mathsf{E}}\Big{[}\tilde{e}_{k+1}\big{|}\mathcal{I}^{e}_{k},\delta_{k}\Big{]}^{T}\operatorname{\mathsf{E}}\Big{[}\tilde{e}_{k+1}\big{|}\mathcal{I}^{e}_{k},\delta_{k}\Big{]}$
$\quad+\operatorname{tr}\operatorname{\mathsf{cov}}\Big{[}\tilde{e}_{k+1}\big{|}\mathcal{I}^{e}_{k},\delta_{k}\Big{]}.$
By Lemmas 4 and 5, we can write
$\tilde{e}_{k+1}$ $=(1-\delta_{k}\gamma_{k})A_{k}\tilde{e}_{k}+K_{k+1}\nu_{k+1}.$
Accordingly, the conditional mean and the conditional covariance of $\tilde{e}_{k+1}$ given $\mathcal{I}^{e}_{k}$ and $\delta_{k}$ are obtained as
$\operatorname{\mathsf{E}}\Big{[}\tilde{e}_{k+1}\big{|}\mathcal{I}^{e}_{k},\delta_{k}\Big{]}$ $=(1-\lambda^{c}\delta_{k})A_{k}\breve{e}_{k}$
$\operatorname{\mathsf{cov}}\Big{[}\tilde{e}_{k+1}\big{|}\mathcal{I}^{e}_{k},\delta_{k}\Big{]}$ $=(\lambda^{c}\delta_{k}-{\lambda^{c}}^{2}\delta_{k})A_{k}\breve{e}_{k}\breve{e}_{k}^{T}A_{k}^{T}$
$\quad+(1-\lambda^{c}\delta_{k})A_{k}R_{k}A_{k}^{T}$
$\quad+K_{k+1}N_{k+1}K_{k+1}^{T}$
where we used the fact that $\nu_{k+1}$ has zero mean and is independent of $\tilde{e}_{k}$ and $\gamma_{k}$, and the properties of covariances. Thus, we get
$\operatorname{\mathsf{E}}\Big{[}\tilde{e}_{k+1}^{T}\tilde{e}_{k+1}\big{|}\mathcal{I}^{e}_{k},\delta_{k}\Big{]}$ $=(1-\lambda^{c}\delta_{k})\breve{e}_{k}^{T}A_{k}^{T}A_{k}\breve{e}_{k}$
$\quad+(1-\lambda^{c}\delta_{k})\operatorname{tr}(A_{k}R_{k}A_{k}^{T})$
$\quad+\operatorname{tr}(K_{k+1}N_{k+1}K_{k+1}^{T}).$
This implies that $\operatorname{\mathsf{E}}[\tilde{e}_{k+1}^{T}\tilde{e}_{k+1}|\mathcal{I}^{e}_{k},\delta_{k}]$ is a symmetric function of $\breve{e}_{k}$ and a function of $R_{k}$, and that
$\chi^{1}_{k}=\lambda^{c}\breve{e}_{k}^{T}A_{k}^{T}A_{k}\breve{e}_{k}+\lambda^{c}\operatorname{tr}(A_{k}R_{k}A_{k}^{T}).$
Therefore, we deduce that $\chi^{1}_{k}$ is also a symmetric function of $\breve{e}_{k}$ and a function of $R_{k}$. Besides, by Lemmas 4 and 5, we can write
$\breve{e}_{k+1}$ $=(1-\bar{\theta}_{k})A_{k}\breve{e}_{k}+K_{k+1}\nu_{k+1}$ (52)
$R_{k+1}$ $=(1-\bar{\theta}_{k})A_{k}R_{k}A_{k}^{T}+\bar{\bar{\theta}}_{k}A_{k}\breve{e}_{k}\breve{e}_{k}^{T}A_{k}^{T}$ (53)
where $\bar{\theta}_{k}=\theta_{k}\delta_{k}\gamma_{k}+\lambda^{c}(1-\theta_{k})\delta_{k}$ and $\bar{\bar{\theta}}_{k}=(1-\theta_{k})(\lambda^{c}\delta_{k}-{\lambda^{c}}^{2}\delta_{k})$. By the hypothesis, $V^{e}_{k+1}(\breve{e}_{k+1},R_{k+1})$ is a symmetric function of $\breve{e}_{k+1}$ and a function of $R_{k+1}$. Plugging (52) and (53) into $V_{k+1}(\breve{e}_{k+1},R_{k+1})$, we can calculate $\operatorname{\mathsf{E}}[V_{k+1}(\breve{e}_{k+1},R_{k+1})|\mathcal{I}^{e}_{k},\delta_{k}]$ while all $\nu_{k+1}$, $\theta_{k}$, and $\gamma_{k}$ are averaged out. A straightforward consequence is that we can express $\operatorname{\mathsf{E}}[V_{k+1}(\breve{e}_{k+1},R_{k+1})|\mathcal{I}^{e}_{k},\delta_{k}]$ as a symmetric function of $\breve{e}_{k}$ and a function of $R_{k}$, since $\nu_{k+1}$ is a Gaussian variable with zero mean. Therefore, we deduce that $\chi^{2}_{k}$ is also a symmetric function of $\breve{e}_{k}$ and a function $R_{k}$. Thus, using (51), we conclude that the claim holds. ∎